What is the minimum value of the function y = 3x2 + 2x – 8?
and
The equation h = 40t – 16t2 describes the height h, in feet, of a ball that is thrown straight up as a function of the time t, in seconds, that the ball has been in the air. At what time does the ball reach its maximum height? What is the maximum height?

Answer by stanbon(75887)

What is the minimum value of the function y = 3x2 + 2x – 8?
Vextex occurs where x = -b/(2a) = -2/(-6) = 1/3
f(1/3) = 3(1/3)^2+2(1/3)-8 = (1/3) + 2/3 -8 = -7
Min at (1/3,-7)
and
The equation h(t) = 40t – 16t2 describes the height h, in feet, of a ball that is thrown straight up as a function of the time t, in seconds, that the ball has been in the air. At what time does the ball reach its maximum height?
t = -b/(2a) = -40/(2*-16) = 5/4 seconds
-----------------------------------------
What is the maximum height?
h(5/4) = 40(5/4)-16(5/4)^2 = 50-25 = 25 ft
================================================
Cheers,
Stan H.
===============